### v7

@@ Abstract
Swarms of autonomous interactive drones can provide compelling sensing capabilities in Smart City applications, such as traffic monitoring. This paper focuses on the task assignment problem for large-scale spatio-temporal sensing by a drone swarm. However, existing approaches have distinct challenges: distributed evolutionary optimization, such as collective learning, lacks long-term adaptability in dynamic environments, while deep reinforcement learning (DRL) is limited to scale effectively due to the curse of dimensionality. Therefore, this paper proposes a novel synergetic optimization approach by integrating long-term DRL and short-term collective learning. Through this approach, each drone independently and proactively determines its flying direction and recharging location using DRL, while evolving their navigation and sensing policies through collective learning based on a structured tree communication model. Extensive experiments with datasets generated from realistic urban mobility demonstrate an outstanding performance of the proposed solution in complex scenarios. New insights show that this approach provides a win-win synthesis of short-term and long-term strategies for drone-based traffic monitoring, with short-term methods addressing training complexity and energy management, while long-term methods preserving high sensing performance.
@@ 1 Introduction
Unmanned Aerial Vehicles (UAVs), referred to as drones, can organize themselves into swarms, fostering collaboration and efficiency in sensor data collection within Smart Cities. With their mobility, autonomy, and diverse sensors, drones have been widely used in several applications, such as monitoring traffic congestion Barmpounakis and Geroliminis (2020), mapping natural disasters Deng et al. (2020) and smart agriculture Liu et al. (2024). Despite these advantages, drones are constrained by limited battery capacity, which impacts their spatio-temporal coverage. This limitation necessitates sophisticated autonomous control systems and precise task assignment within drone swarms. This involves the assignment of different sensing tasks to each drone while meeting the sensing requirements, drone capabilities and constraints Poudel and Moh (2022). Recently, scholars have solved the NP-hard combinatorial optimization problem of task assignment for spatio-temporal sensing, using evolutionary algorithms ranging from genetic algorithms to swarm intelligence methods Poudel and Moh (2022); Zhou et al. (2020). To enhance robustness and minimize the effects of individual drone failures, distributed optimization approaches have been introduced. These approaches allow drones to autonomously self-organize and self-assign tasks in a decentralized manner, preserving the autonomy of individual drones. However, these distributed approaches are reactive to the current state of the environment rather than proactive to address long-term operational challenges strategically.
Strategic foresight in navigation and sensing is crucial for optimizing drone operations in dynamic environments, significantly influenced by flying directions and future sensor data requirements. For example, drones aware of an expected increase in traffic can proactively fly to those areas, improving tasks such as vehicle detection, even if the short-term benefits are not evident. This "slower is faster" effect emphasizes the importance of long-term planning for overall effectiveness in drone sensing. To address long-term optimization, deep reinforcement learning (DRL) has been introduced, utilizing the Bellman equation to consider cumulative rewards Cui et al. (2019). DRL based on deep neural networks support drones, as heterogeneous agents, to manage complex state spaces and adapt to changing environments.
However, traditional DRL approaches cannot scale well with a large number of agents and extended time spans due to the curse of dimensionality. With more agents, coordinating them becomes exponentially more complex, leading to computational challenges. Furthermore, DRL methods need to account of environmental uncertainties and the long-term consequences of agents’ actions, which become increasingly complex as the time horizon extends. For example, drones often make flight path decisions over short intervals, such as seconds or minutes, but face significant challenges when required to adapt their policies for a long duration (e.g., a full day). Meanwhile, the extended operation is longer than the maximum flight time of drones, enabling them to return to a base station for recharging and depart from that place for the subsequent task. The selection of these recharging places constraints the drones’ flight ranges, which in turn affects their sensing performance.
Therefore, in this paper, we address a novel task assignment problem for large-scale spatio-temporal sensing, focusing on the navigation (determining flight path), sensing (scheduling data collection) and recharging (choosing charging stations as destinations) of drone swarms in changing environments. To jointly manage these tasks, we propose a new synergetic optimization approach to coordinate drone swarms through Distributed Optimization and (deep) Reinforcement Learning (DO-RL), which involves a multi-agent DRL algorithm that builds upon on an iterative collective learning approach Pournaras et al. (2018); Pournaras (2020). It allows drones to determine long-term flying directions from departures to destinations (i.e., charging stations) via DRL, while autonomously evolving their short-term navigation and sensing policies based on shared plans from other drones through collective learning. The distributed sensing results generated during operations serve as rewards and observations for further DRL training. These results are aggregated and shared using a structured tree communication model, where each drone interacts with its parent and child nodes, making independent decisions. This ensures scalability (supporting a large number of agents), efficiency (low communication and computational cost), decentralization and resilience Majumdar et al. (2023); Pournaras et al. (2020); Nikolic and Pournaras (2019). By employing this collaborative approach, collective learning reduces the frequency of decision-making and mitigates inefficient actions, effectively compensating for the limitations of DRL while retaining its strengths in long-term planning and optimization. In addition, a set of auxiliary approaches is introduced, including plan generation, plan selection, periodic state updates and a centralized training and decentralized execution module. To validate DO-RL, this paper conducts experimental evaluations with real-world data and emulated traffic monitoring scenarios, comparing it against state-of-the-art baseline methods.
@@
The contributions of this paper are outlined as follows: (i) The first study of task assignment for large-scale, dynamic spatio-temporal sensing by a swarm of drones, concentrating on optimizing the flight paths, data collection strategies, and recharging. (ii) A novel synergetic optimization approach, DO-RL, that tackles this problem by integrating both multi-agent DRL and distributed optimization based on a structured tree communication model. (iii) A testbed for extensive experimentation with realistic traffic patterns and real-world transport networks. (iv) The validation for the superior performance of the proposed approach over three state-of–the-art methods. (v) An open-source implementation of the proposed approach and an open dataset [^1] https://doi.org/10.6084/m9.figshare.24476533.v2 containing all plans of the studied scenario. They can be used as benchmarks to encourage further research on this problem.
@@
Alighanbari and How (2006)
Chen et al. (2022)
Qin et al. (2022)
Ding et al. (2021)
Zhao et al. (2022)
Li et al. (2023)
Samir et al. (2020)
@@
Long-term efficiency
Scalability
Resilience
Adaptability
2 Related Work
Sensing task assignment. The UAV task assignment problem for spatio-temporal sensing has been traditionally defined as the Traveling Salesmen Problem Bartolini et al. (2019). This involves finding the optimal assignment of tasks to drones at specific locations, while considering constraints such as task urgency, time scheduling, and flying costs. The evolutionary optimization methods, such as particle swarm optimization Phung and Ha (2021), and genetic algorithm Pehlivanoglu and Pehlivanoglu (2021), have been introduced to formulate the problem of reaching optimal sensing efficiency in terms of coverage, events detection rate and the cost of flight paths. They rely on adaptive exploration of the solution space to optimize task allocation based on shared information and local interactions. However, these methods are centralized, wherein the failure of a central control station can lead to several system disruptions without autonomous recovery mechanisms Omidshafiei et al. (2017).
Distributed optimization. Several distributed task assignment algorithms have been proposed. One such algorithm is the robust decentralized task assignment (RDTA) Alighanbari and How (2006) that enhances robustness and reduces communication costs by employing decentralized planning for drones. Another one is the consensus-based bundle algorithm (CBBA) Chen et al. (2022), which combines market-based mechanisms and situation awareness to converge and avoid task conflicts among drones. Moreover, a planning-based coordination method using collective learning, named Economic Planning and Optimized Selections (EPOS), enables drones to optimize their navigation and sensing options under battery constraints Qin et al. (2022); Qin and Pournaras (2023). However, long-term sensing scenarios pose challenges. Drones tend to prioritize immediate maximum returns in the short term, often neglecting the need for strategic decision-making for overall benefits.
DRL-based drone sensing. The deep reinforcement learning (DRL) algorithm proves to be effective for addressing distributed sensing with drones, managing complex optimization goals, energy constraints, and numerous heterogeneous agents. Ding et al. Ding et al. (2021) integrated mobile crowdsensing (i.e., human equipped with mobile sensing devices) into UAV sensing, overcoming the limitations of both human mobility and drone battery life. Omoniwa et al. Omoniwa et al. (2023) presented a decentralized DRL approach that improves drone energy efficiency in providing wireless connectivity by sharing information with neighbouring drones. Zhao et al. Zhao et al. (2022) proposed a comprehensive solution for sustainable urban-scale sensing in open parking spaces. Their solution incorporates the task selection and scheduling using DRL, and adaptive recharging assignment. However, these approaches lack autonomy for drones in self-assignment tasks, reducing adaptability to the real-time environmental changes and increasing vulnerability to agent failures Omidshafiei et al. (2017). Additionally, they are confined to a small-scale task assignment problem due to the training complexity.
Evolutionary algorithm and DRL. In this optimization process, the related synergetic optimization algorithms integrate the full learning and optimization capabilities of evolutionary algorithms and DRL. They independently solve sub-problems to generate partial solutions, which are then combined to form a complete solution Li et al. (2024). Earlier work Majumdar et al. (2020); Aydeniz et al. (2023) has proposed optimizing team policies using evolutionary algorithms with team rewards, while simultaneously optimizing individual policies through DRL with agent rewards, thus bypassing the non-stationary problem in multi-agent DRL. Li et al. Li et al. (2023) introduce a hybrid framework using shared observation encoders and independent linear policy representations to enhance collaboration on complex task, ensuring stable evolution through agent-level crossover and mutation. While benchmarks, such as MuJoCo Peng et al. (2021) and Starcraft II Samvelyan et al. (2019), are useful for evaluation of these approaches, these benchmarks may struggle to adapt to the continuously evolving and unanticipated dynamics typical of real-world scenarios, such as fluctuating traffic patterns in drone-based monitoring systems.
Transportation application. In the application of transportation system, drones rely on accurate vehicle observation to detect early signs of traffic congestion. This enables traffic operators to apply mitigation measures, reducing the carbon footprint in one of the world’s highest-emission sectors. Bakirci et alBakirci (2024) explored vehicle detection using the YOLOv8 algorithm on aerial images captured by a custom UAV. Samir et al. Samir et al. (2020) leveraged a DRL algorithm to optimize the trajectories of UAVs, with the objective of minimizing the age of information of collected data. While these approaches demonstrate high performance in the detection of traffic vehicles, scalability becomes a challenge since they only deal with a small number of drones within a short duration.
@@ 2 Related Work
In summary, previous approaches have either neglected the long-term advantages of strategic sensing, or failed to effectively adapt to failures and scale without encountering centralized bottlenecks. To overcome these shortcomings, the proposed approach empowers drones to make informed and proactive decisions based on predicted traffic flow, while enabling them to independently adjust to changing conditions and unexpected events, ensuring both scalability and resilience without relying on centralized control.

### v8

@@ Abstract
This paper addresses the problem of autonomous task allocation by a swarm of autonomous, interactive drones in large-scale, dynamic spatio-temporal environments. When each drone independently determines navigation, sensing, and recharging options to choose from such that system-wide sensing requirements are met, the collective decision-making becomes an NP-hard decentralized combinatorial optimization problem. Existing solutions face significant limitations: distributed optimization methods such as collective learning often lack long-term adaptability, while centralized deep reinforcement learning (DRL) suffers from high computational complexity, scalability and privacy concerns. To overcome these challenges, we propose a novel hybrid optimization approach that combines long-term DRL with short-term collective learning. In this approach, each drone uses DRL methods to proactively determine high-level strategies, such as flight direction and recharging behavior, while leveraging collective learning to coordinate short-term sensing and navigation tasks with other drones in a decentralized manner. Extensive experiments using datasets derived from realistic urban mobility demonstrate that the proposed solution outperforms standalone state-of-the-art collective learning and DRL approaches by $27.83\%$ and $23.17\%$ respectively. Our findings highlight the complementary strengths of short-term and long-term decision-making, enabling energy-efficient, accurate, and sustainable traffic monitoring through swarms of drones.
@@ 1 Introduction
Unmanned Aerial Vehicles (UAVs), referred to as drones, can organize themselves into swarms, fostering collaboration and efficiency in sensor data collection within Smart Cities. With their mobility, autonomy, and intelligence, drones have been widely used in several applications, such as monitoring traffic congestion barmpounakis2020new, mapping natural disasters deng2020two and smart agriculture liu2024complex. Despite these advantages, drones are constrained by limited battery capacity, which impacts their spatio-temporal coverage. This limitation necessitates sophisticated autonomous control systems and precise task allocation within drone swarms. This involves the allocation of different sensing tasks to each drone while meeting the sensing requirements, drone capabilities and constraints poudel2022task. Recently, scholars have solved the NP-hard combinatorial optimization problem of assigned sensing tasks by drone swarms, using task allocation algorithms ranging from genetic algorithms to swarm intelligence methods poudel2022task; zhou2020uav.
To enhance robustness and minimize the effects of individual drone failures, distributed optimization approaches, such as decentralized collective learning, have been introduced pournaras2018decentralized. These approaches coordinate agents to autonomously self-organize and self-assign task plans in a decentralized manner, respecting the autonomy and operational flexibility of individual agents. Due to their scalability (supporting a large number of agents), efficiency (low communication and computational cost), decentralization and resilience pournaras2018decentralized, they are well-suited for tackling the autonomous task allocation problem for large-scale spatio-temporal sensing (i.e., a decentralized combinatorial optimization problem), which is the focus of this paper. In particular, they solve the navigation (determining flight path), sensing (scheduling data collection) and recharging (choosing charging stations as destinations) of drone swarms. This approach is especially valuable in scenarios involving multi-modal sensing requirements, where heterogeneous drones, equipped with diverse sensors (e.g., camera, LiDAR, thermal), collaborate across different institutions to perform coordinated sensing while preserving the privacy of individual agents.
However, distributed optimization approaches are reactive to the current state of the environment rather than proactive to address strategically long-term operational challenges. When the time horizon for autonomous task allocation by drone swarms extends-from short durations constrained by flight range to longer periods spanning a whole day-the sensing environment (e.g., traffic flow) becomes increasingly dynamic and unpredictable. To remain effective in such evolving conditions, drone swarms must incorporate strategic foresight into their decision-making, i.e., anticipating flying directions and future sensor data collective requirements. For example, drones aware of an expected increase in traffic can proactively fly to those areas, even if those areas are currently sparsely populated with vehicles. They can also land on base stations near those areas for recharging and executing sensing tasks in the next time period. As a result, the strategic decision-making improves tasks of vehicle detection, even if the short-term benefits are not evident. This "slower is faster" effect emphasizes the importance of long-term planning for overall effectiveness in drone sensing.
To address long-term optimization, deep reinforcement learning (DRL) has been introduced, utilizing the Bellman equation to consider cumulative rewards cui2019multi. DRL based on deep neural networks support drones, as heterogeneous agents, to manage complex state spaces and adapt to changing environments cui2019multi. Nevertheless, traditional DRL approaches struggle with the autonomous task allocation problem because of their inherent scalability limitations, high complexity, and reliance on centralized control. As sensing tasks prolong to an entire day, the number of decisions per drone increases significantly since DRL typically operates on short time intervals (e.g., seconds or minutes). This leads to an exponential growth of the action space, increasing training complexity and hindering effective exploration of optimal action to allocate sensing tasks. The curse of dimensionality could be furt her exacerbated by a larger number of agents/drones. Moreover, centralized training requires access to private or sensitive information of drones, including battery level and collected sensor data, thereby compromising the autonomy and privacy of heterogeneous drones. Notably, these limitations of DRL are where distributed optimization methods excel, offering scalable, decentralized, and privacy-preserving solutions, motivating the study of a hybrid approach.
Therefore, in this paper, we propose a new hybrid optimization approach to coordinate drone swarms through Distributed Optimization and (deep) Reinforcement Learning (DO-RL), which involves a multi-agent DRL algorithm that builds upon on a decentralized collective learning approach pournaras2018decentralized. This approach leverages a hierarchical decision-making framework that uses both high-level long-term policies and low-level short-term coordination. Specifically, DO-RL allows drones to determine flying directions from departures to destinations (i.e., charging stations) over multiple time periods via DRL, while autonomously selecting their navigation and sensing plans within a time period through collective learning. The plans are generated locally and autonomously by drones without leaking their private data to others. As a consequence, the low-level collective learning not only reduces the number of actions in high-level DRL and training complexity, but efficiently coordinates each drone to select plans and observe the aggregated plans of others through a tree communication topology. The aggregated plans are used to calculate the reward of DRL for optimizing sensing performance.
@@
The contributions of this paper are outlined as follows: (i) The first study of autonomous task allocation for large-scale, dynamic spatio-temporal sensing by a swarm of drones, concentrating on optimizing the flight paths, data collection strategies, and recharging. (ii) A novel hybrid optimization approach, DO-RL, that tackles this problem by integrating both multi-agent DRL and decentralized collective learning. It includes a set of auxiliary methods such as DRL-based scheduling, local plan generation, iterative plan selection, periodic state update, actor-critic networks, and policy proximal optimization. (iii) A testbed for extensive experimentation with realistic traffic patterns and real-world transport networks to validate the higher energy-efficient and accurate vehicle observation of DO-RL compared to state-of-the-art baseline methods. (iv) Quantitative findings that show how the proposed hybrid approach achieves a win-win synthesis of long-term learning and short-term optimization in both sensing and charging under a large spectrum of factors such as drone density, number of time periods, and vehicle density. (v) An open-source implementation of the proposed approach and an open dataset [^1] Available at: https://doi.org/10.6084/m9.figshare.24476533.v2 containing all plans of the studied scenario. They can be used as benchmarks to encourage further research on this problem.
2 Related Work
The sensing task allocation problem of drones has been traditionally defined as the traveling salesmen problem or vehicle routing problem, which is a NP-hard combinatorial optimization problem bartolini2019task. This involves finding the optimal assignment of tasks to drones at specific locations, while considering constraints such as task urgency, time scheduling, and flying costs. The task allocation algorithms, such as Particle Swarm Optimization (PSO) phung2021safety, and genetic algorithm pehlivanoglu2021enhanced, have been introduced to formulate the problem of reaching optimal sensing efficiency in terms of coverage, events detection rate and the cost of flight paths. They rely on adaptive exploration of the solution space to optimize task allocation based on shared information and local interactions. However, these methods are centralized, wherein the failure of a central control station can lead to several system disruptions without autonomous recovery mechanisms liu2023decentralized.
To solve the decentralized combinatorial optimization problem, several distributed task allocation algorithms have been proposed: One such algorithm is the Robust Decentralized task allocation (RDTA) alighanbari2006robust that enhances robustness and reduces communication costs by employing decentralized planning for drones. Another one is the Consensus-Based Bundle Algorithm (CBBA) bertuccelli2009real, which combines market-based mechanisms and situation awareness to converge and avoid task conflicts among drones. The COHDA hinrichs2013cohda optimization generalizes well in different communication structures among drones with full view of the system, but requires each drone to exchange information with all of others. Moreover, a decentralized multi-agent collective learning approach, named Economic Planning and Optimized Selections (EPOS), allows agents to exchange information over a tree-structured communication topology, achieving low computational and communication overhead pournaras2018decentralized. Compared to the relevant algorithm of COHDA, this approach has lower cost of message exchange and higher cost-effectiveness in terms of optimality pournaras2018decentralized. Moreover, EPOS enables each drone to share the aggregated plans of navigation and sensing to others, without exposing the private information of drone specifications, battery and sensor data qin2024m. Nevertheless, long-term sensing scenarios pose challenges. Drones tend to prioritize immediate maximum returns in the short term, often neglecting the need for strategic decision-making for long-term benefits.
The deep reinforcement learning (DRL) algorithm proves to be effective for addressing the combinatorial optimization problem in the evolving multi-drone sensing environments, by transforming the problem in to a sequential decision-making process barrett2020exploratory; mazyavkina2021reinforcement. Ding et al. ding2021crowdsourcing integrate mobile crowdsensing (i.e., human equipped with mobile sensing devices) into UAV sensing, overcoming the limitations of both human mobility and drone battery life. They use the centralized training decentralized execution framework where drones can share auxiliary information during training but act independently during execution, preserving decentralization. Zhao et al. zhao2022dronesense propose a comprehensive solution for sustainable urban-scale sensing in open parking spaces. Their solution incorporates the task selection and scheduling using DRL, and adaptive recharging assignment. Omoniwa et al. omoniwa2023communication present a multi-agent decentralized DRL approach that improves drone energy efficiency in providing wireless connectivity by sharing information with neighboring drones. However, these approaches are confined to a small-scale task allocation problem due to high training complexity.
The scalability of multi-agent DRL has been widely studied. Previous work combines DRL with mean-field methods yang2018mean; chen2020mean that transforms the complex interactions within the population of agents into a simplified interaction between a single agent and the average effect from the overall population. Other work focus on reducing the joint action-state space to improve the exploration efficiency kanervisto2020action; majeed2021exact; machado2023temporal. Apart from the action abstraction approaches that rely on predefined and offline abstractions, such as masking, sequentialization and temporal abstraction, the hierarchical reinforcement learning jendoubi2023multi; xu2023haven has been introduced to provide flexibility in handling non-linear and complex task environments. It structures decision-making into high- and low-level tasks, enabling agents to learn complex behaviors more efficiently by decomposing the problem. Furthermore, a recent work by Qiu et al. qiu2022dimes propose DIMES, a differentiable meta-solver that leverages the compact continuous space and DRL to address traveling salesmen problems at large scales (up to 10000 nodes in a graph). Nevertheless, these approaches lack autonomy for drones in autonomous allocation tasks, increasing vulnerability to agent failures and private information leakage.
In the application of transportation system, drones rely on accurate vehicle observation to detect early signs of traffic congestion. This enables traffic operators to apply mitigation measures, reducing the carbon footprint in one of the world’s highest-emission sectors. Samir et al. samir2020age leverage a DRL algorithm to optimize the trajectories of UAVs, with the objective of minimizing the age of information of collected data. Bakirci et albakirci2024enhancing explore vehicle detection using the YOLOv8 algorithm on aerial images captured by a custom UAV. While these approaches demonstrate high performance in the detection of traffic vehicles, scalability becomes a challenge since they only deal with a small number of drones within a short duration.
@@ 2 Related Work
bertuccelli2009real
hinrichs2013cohda
qin2024m
ding2021crowdsourcing
zhao2022dronesense
qiu2022dimes
samir2020age
@@
Long-term adaptability
Scalability with low complexity
Autonomy and privacy-preservation
@@
In summary, previous approaches have either neglected the long-term advantages of strategic sensing, or failed to scale without encountering centralized bottlenecks. To overcome these shortcomings, the proposed approach empowers drones to make informed and proactive decisions based on predicted traffic flow, while enabling them to independently adjust to changing conditions and unexpected events, ensuring both scalability and adaptability without relying on centralized control. Table 1 illustrates the comparison to related work in different types of criteria.